OpenStudy (anonymous):

what can you conclude if a and b are nonzero integers such that a divides b and b divides a? please helppp

OpenStudy (wwwhhhaaattt?):

What do you think it is?

OpenStudy (anonymous):

its our homework in number theory and i cant anwer this question... its all about divisibility

OpenStudy (wwwhhhaaattt?):

Do you know your Times tables?

OpenStudy (anonymous):

yeah i knoww.

OpenStudy (wwwhhhaaattt?):

Ok, Then I want to see you try.

OpenStudy (anonymous):

try? how?

OpenStudy (wwwhhhaaattt?):

Try.

OpenStudy (anonymous):

like all of em?

jimthompson5910 (jim_thompson5910):

a divides b means b = a*k for some integer k (basically 'a' is a factor of 'b') b divides a means a = b*m for some integer m (basically 'b' is a factor of 'a')

OpenStudy (wwwhhhaaattt?):

Just try

jimthompson5910 (jim_thompson5910):

so use the two equations b = a*k a = b*m to help answer the question

jimthompson5910 (jim_thompson5910):

does that help @alesisaminion ?

OpenStudy (anonymous):

@jim_thompson5910 yeah :)) thanks a lot :)xx

jimthompson5910 (jim_thompson5910):

ok great, I'm glad it does

jimthompson5910 (jim_thompson5910):

tell me what you can conclude about m and k

OpenStudy (anonymous):

@jim_thompson5910 look at your message

jimthompson5910 (jim_thompson5910):

you should have something along these lines in your steps/work b = a*k b = b*m*k ... replace 'a' with 'b*m' b = b*(m*k) b/b = [b*(m*k)]/b ... divide both sides by b 1 = m*k m*k = 1 So I started with one equation, b = a*k, and ended up with m*k = 1. If m*k = 1, then what must m and k be? Hint: m and k are integers (they can be the same integer)